### 10 Mar 2026 v2

# Joint distribution of leftmost digits in positional notation and Schanuels’s conjecture

Abstract
Assume that $n\geq 2$ and $B=(b_{1},...,b_{n})$ has distince integer entries $\geq 3.$ For $x>0$ let $d_{B}(x):=(d_{b_{1}}(x),...,d_{b_{n}}(x))$ where $d_{b_{i}}(x)\in\{1,...,b_{i}-1\}$ is the leftmost digit in the base-$b_{i}$ positional notation representation of $x.$ We prove that if $d_{B}$ is surjective, then $\ln b_{i}$ and $\ln b_{j}$ are rationally independent whenever $i\neq j.$ We prove the converse for $n=2,$ and for $n\geq 3$ if $\{\ln p:p\mbox{ prime}\}$ is algebraically independent, a condition implied by Schanuel’s conjecture about transcendental numbers.
2010 Mathematics Subject Classification: 11A63 11J81 37A17  [^1] This work is supported by the Krasnoyarsk Mathematical Center and financed by the Ministry of Science and Higher Education of the Russian Federation (Agreement No. 075-02-2026-1314).
1 Positional notation
$\mathbb{Z},\mathbb{Q},\mathbb{R},\mathbb{R}_{+}$ denote integer, rational, real and positive numers, and iff mean if and only if. For integers $b\geq 3,$ $e\geq 1,$ $j\in\{1,...,b-1\}$ define $d_{b}:\mathbb{R}_{+}\mapsto\{1,...,b-1\}$ and $S(b,e,j)\subset\{1,...,b^{e}-1\}$ by
$d_{b}(x)=j\iff\log_{b}(x)\in[\log_{b}(j),\log_{b}(j+1))+\mathbb{Z},\ \ x\in\mathbb{R}_{+},$ (1)
$S(b,e,j):=\bigcup_{\ell=0}^{e-1}\left(b^{\ell}j+\{0,...,b^{\ell}-1\}\right).$ (2)
$d_{b}(x)$ is the leftmost ’digit’ in the base-$b$ positional notation representation of $x.$
Lemma 1
$d_{b^{e}}(x)$ determines $d_{b}(x)$ by
$d_{b}(x)=j\iff d_{b^{e}}(x)\in S(b,e,j).$ (3)
Proof 1
The identity $\log_{b}(x)=e\,\log_{b^{e}}(x)$ implies
$d_{b}(x)=j\iff\log_{b^{e}}(x)\in[\log_{b^{e}}(j),\log_{b^{e}}(j+1))+e^{-1}\mathbb{Z}$
$\iff\log_{b^{e}}(x)\in[\log_{b^{e}}(j),\log_{b^{e}}(j+1))+e^{-1}\{0,...,e-1\}+\mathbb{Z}$
$\iff\log_{b^{e}}(x)\in\bigcup_{\ell=0}^{e-1}[\log_{b^{e}}(b^{\ell}j),\log_{b^{e}}(b^{\ell}(j+1)))+\mathbb{Z}$
$\iff d_{b^{e}}(x)\in\bigcup_{\ell=0}^{e-1}\mathbb{Z}\cap[b^{\ell}j,b^{\ell}(j+1))\iff d_{b^{e}}(x)\in S(b,e,j).$
2 Joint distribution for two bases
Assume that $B=(b_{1},b_{2})$ has distinct integer entrues $\geq 3.$ Define
$d_{B}:\mathbb{R}_{+}\mapsto\{1,...,b_{1}-1\}\times\{1,...,b_{2}-1\}$
by $d_{B}(x):=(d_{b_{1}}(x),d_{b_{2}}(x)).$ $\ln b_{1}$ and $\ln b_{2}$ are rationally dependent iff there exist integers $a,e_{1},e_{2}$ such that $a\geq 2,$ $gcd(e_{1},e_{2})=1,$ and $b_{i}=a^{e_{i}},i=1,2.$ Then $b:=b_{1}^{e_{2}}=b_{2}^{e_{1}}$ and Lemma 1 implies
$d_{B}(x)=(j_{1},j_{2})\iff d_{b}(x)\in S(b_{1},e_{2},j_{1})\cap S(b_{2},e_{1},j_{2}).$ (4)
Example 1
If $B=(4,8),$ then $a=2,e_{1}=2,e_{2}=2$ and $b=64.$ We used (4) to compute $(j_{1},j_{2})=d_{B}(x)$ as a function of $d_{64}(x)\in\{1,...,63\}$ displayed in the table below. $d_{B}$ is not surjective because its image excludes the six pairs $(2,3),(2,6),(2,7),(3,2),(3,4),(3,5).$
$\begin{array}[]{cccccccc}&&j_{1}=1&&j_{1}=2&&j_{1}=3\\ j_{2}=1&&1&&8-11&&12-15\\ j_{2}=2&&16-23&&2&&\emptyset\\ j_{2}=3&&24-32&&\emptyset&&3\\ j_{2}=4&&4&&32-39&&\emptyset\\ j_{2}=5&&5&&40-47&&\emptyset\\ j_{2}=6&&6&&\emptyset&&48-55\\ j_{2}=7&&7&&\emptyset&&56-63\end{array}$
Define the circle group $\mathbb{T}:=\mathbb{R}/\mathbb{Z}$ and $n$-dimensional torus group $\mathbb{T}^{n}=\mathbb{R}^{n}/\mathbb{Z}^{n}.$ For $\omega=(\omega_{1},...,\omega_{n})\in\mathbb{R}^{n}$ define the homomorphism $\varphi_{\omega}:\mathbb{R}\mapsto\mathbb{T}^{n}$ by $\varphi_{\omega}(t):=t\omega+\mathbb{Z}^{n}.$
Lemma 2
The image of $\varphi_{\omega}$ is dense in $\mathbb{T}^{n}$ iff $\omega_{1},...,\omega_{n}$ are rationally indpendent.
Proof 2
This landmark result was proved in 1884 by Leopold Kronecker [2]. A modern English langauge proof is given in Proposition 1.5.1 in [1].
Theorem 1
If $\ln b_{1}$ and $\ln b_{2}$ are rationally dependent and integers $a,e_{1},e_{2}$ satisfy $a\geq 2,$ $gcd(e_{1},e_{2})=1,$ and $b_{i}=a^{e_{i}}\geq 3,i=1,2,$ then $(j_{1},j_{2})\in\{1,...,b_{1}-1\}\times\{1,...,b_{2}-1\}$ is in the image of $d_{B}$ iff there exists an integer $c$ such that
$\frac{j_{1}}{j_{2}+1}<a^{c}<\frac{j_{1}+1}{j_{2}}.$ (5)
Since $a^{c}$ never belongs to the open interval $(\frac{1}{2},1),$ $d_{B}$ is not surjective since $(2,3)$ is not contained in its image. If $\ln b_{1}$ and $\ln b_{2}$ are rationally independent, then $d_{B}$ is surjective.
Proof 3
If $\ln b_{1}$ and $\ln b_{2}$ are rationally dependent and $x\in\mathbb{R}_{+}$ then $d_{B}(x)=(j_{1},j_{2})$ iff there exist integers $d_{1}$ and $d_{2}$ such that
$x\in a^{e_{1}d_{1}}[j_{1},j_{1}+1)\cap a^{e_{2}d_{2}}[j_{2},j_{2}+1).$ (6)
Then (5) holds for $c=e_{2}d_{2}-e_{1}d_{1}.$ Conversly, if (5) holds then since $gcd(e_{1},e_{2})=1,$ there exist integers $d_{1}$ and $d_{2}$ such that $c=e_{2}d_{2}-e_{1}d_{1}$ and then (6) holds. If $\ln b_{1}$ and $\ln b_{2}$ are rationally independent. Then $\omega_{1}:=\frac{1}{\ln b_{1}}$ and $\omega_{2}:=\frac{1}{\ln b_{2}}$ are rationally independent so Lemma 2 and the surjectivtivity of $\ln:\mathbb{R}_{+}\mapsto\mathbb{R}$ implies that the image of the composition $h:=\varphi_{\omega}\circ\ln:\mathbb{R}_{+}\mapsto\mathbb{T}^{2}$ is dense in $\mathbb{T}^{2}.$ Partition $\mathbb{T}^{2}$ into sets
$R(j_{1},j_{2}):=[\log_{b_{1}}(j_{1}),\log_{b_{1}}(j_{1}+1))\times[\log_{b_{2}}(j_{2}),\log_{b_{2}}(j_{2}+1))+\mathbb{Z}^{2},\}.$
[…]
3 Schanuel’s conjecture and its consequences
Schanuel’s conjecture ([3], p. 30–32) asserts: If $c_{1},...,c_{m}$ are rationaly independent complex numbers, then the field $\mathbb{Q}(c_{1},...,c_{m},exp(c_{1}),...,exp(c_{m}))$ has transcendence degree at least $m$ over $\mathbb{Q}.$
Let $\mathbb{Q}(z_{1},...,z_{m})$ denote the field of rational functions in variables $z_{1},...,z_{m}$ represented by the field of meromorphic functions on the $n$-fold product of Riemann spheres $(\mathbb{C}\cup\{\infty\})^{m}.$ Assume that $m\geq 1$ and $p_{1},...,p_{m}$ are distinct primes,
Conjecture 1
$\ln p_{1},...,\ln p_{m}$ are algebraically independent. Equivalently, the homomorphism $\psi:\mathbb{Q}(z_{1},...,z_{m})\mapsto\mathbb{Q}(\ln p_{1},...,\ln p_{m})$ satisfying $\psi(z_{j})=\ln p_{j}$ is an isomorphism.
Lemma 3
Schanuel’s conjecture implies conjecture 1
Proof 4
The unique factorization of positive integers as products of power of primes implies that $\ln p_{1},...,\ln p_{m}$ are rationally independent. Conjecture 1 then follows form Schanuel’s conjecture by defining $c_{i}:=\ln p_{i},i=1,...,n.$
Conjecture 2
If $b_{1},...,b_{n}$ are integers $\geq 3$ such that no pair of numbers in the set $\{\ln b_{1},...,\ln b_{n}\}$ are rationaly dependent, then $\frac{1}{\ln b_{1}},...,\frac{1}{\ln b_{n}}$ are rationally independent.
Proposition 1
Conjecture 1 implies conjecture 2.
Proof 5
Let $p_{1},...,p_{m}$ be the distinct primes that divide $b_{1}\cdots b_{n}.$ Then there exist integers $e_{i,j}\geq 0,i=1,...,n;j=1,...,m$ such that
$b_{i}=p_{1}^{e_{i,1}}\cdots p_{m}^{e_{i,m}},\ \ i=1,...,n.$ (7)
Conjecture 1 implies that $\psi$ is an isomorphism hence $\psi^{-1}(\ln b_{i})=\sum_{j=1}^{m}e_{i,j}z_{j},$ so
$\zeta_{i}:=\psi^{-1}\left(\frac{1}{\ln b_{i}}\right)=\left(\sum_{j=1}^{m}e_{i,j}z_{j}\right)^{-1},\ \ i=1,...,n.$ (8)
$\zeta_{i}(z_{1},...,z_{m})=\infty$ on an (m1)(m-1)-dimensional $\mathbb{C}$-subspaces $S_{i}$ of $\mathbb{C}^{m}\subset(\mathbb{C}\cup\{\infty\})^{m}.$ The hypothesis in conjecture 2 implies that $S_{1},...,S_{m}$ are distinct. Therefore $\zeta_{1},...,\zeta_{m}$ and hence $\frac{1}{\ln b_{1}},...,\frac{1}{\ln b_{1}}$ are rationally independent.
4 Joint distribution for $n\geq 3$ bases
Letf $B=(b_{1},...,b_{n})$ with distinct integer entries $\geq 3$ and defin $\omega:=(\omega_{1},...,\omega_{n}),\omega_{i}:=\frac{1}{\ln b_{i}}.$ Then $d_{B}:\mathbb{R}_{+}\mapsto\{1,...,b_{1}-1\}\times\cdots\times\{1,...,b_{n}-1\}.$
Theorem 2
If some pair of numbers in the set $\{\ln b_{1},...,\ln b_{n}\}$ are rationaly dependent, then $d_{B}$ is not surjective. If no pair of numbers in the set $\{\ln b_{1},...,\ln b_{n}\}$ are rationaly dependent, then conjecture 2, and hence Schanuel’s conjecture, implies $d_{B}$ is surjective.
Proof 6
The first assertion follows from Theorem 1. Assume that no pair of numbers in the set $\{\ln b_{1},...,\ln b_{n}\}$ are rationaly dependent and assume conjecture 2. Then the entries of $\omega$ are rationally independent. Therefore Lemma 2 implies that the image of $h:=\psi_{\omega}\circ\ln:\mathbb{R}_{+}\mapsto\mathbb{T}^{n}$ is dense in $\mathbb{T}^{n}.$ The same argument used in the proof of Theorem 1 shows that $d_{B}$ is surjective.